OpenStudy (anonymous):

Solve the system of equations graphically and check solutions algebraically. How do I start, help!? x+2y=7 3x+y=6

OpenStudy (anonymous):

You would need to graph the two points and find where they intersect (the x and y coordinates) Algebraically: (x + 2y = 7)*1 to x+2y = 7 (3x + y = 6)*2 to 6x +2y = 12 5x = 5 (Subtract the two equations) x = 1 y = 3

OpenStudy (anonymous):

How do you graph that?

OpenStudy (anonymous):

Punch it into a grapher (calculator or online)

OpenStudy (anonymous):

I know how to get to the solution by doing what you did, but how do I solve the system of equations graphically...I don't understand how to punch it in...